Circular No. 8501

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2005 Further to IAUC 8500, S. Nakano forwards the following precise po-

sition from K. Kadota (Ageo, Japan, 0.25-m f/5.0 reflector) from an unfiltered CCD image taken on Mar. 29.737 UT, with the nova at mag 8.2: α = 18^{h}16^{m}58^{s}.96, δ = −25^{o}56′38′′.9 (equinox 2000.0). Nakano also forwards the following position end figures from an unfiltered image taken by S. Wakuda (Yuto-cho, Hamana-gun, Shizuoka-ken, Japan, 0.25-m f/5.0 Schmidt-Cassegrain reflector) on Mar. 29.755, with the nova at mag 7.8: 58^{s}.97, 39′′.1.

L. Kiss, University of Sydney; and A. Derekas, University of New South Wales, write that an optical spectrum (0.05 nm/pixel; range 580– 675 nm), taken with the Australian National University 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring Observatory on Mar. 29.81 UT, shows that the new object announced on IAUC 8500 is indeed a nova. The prominent Hα emission line shows a strong P-Cyg profile, with the main absorption component at −1300 km/s; there might also be a weak second component at −2100 km/s. The FWZI of the Hα line exceeds 5000 km/s. In addition to Hα, a few broad features are visible between 618 and 640 nm, also with P-Cyg-like profiles. The Na D doublet is saturated, indicating high interstellar reddening.

K. Ayani and Y. Kawabata, Bisei Astronomical Observatory (BAO), report that their low-dispersion spectra (range 453–680 nm; resolution 0.6 nm at Hα) of this new object (IAUC 8500), obtained on Mar. 29.8 UT at the BAO 1.01-m telescope, also show it to be a nova, with Hα, Hβ, and Fe II (multiplet 42) lines having P-Cyg profiles. Preliminary analysis of the Hα emission (without decomposition of the absorption component) yields FWHM = 1300 km/s, the minimum of the Hα P-Cyg absorption being blueshifted by 1400 km/s with respect to the emission peak.

SUPERNOVA 2005av IN NGC 6943 M. Salvo and B. Schmidt, Australian National University (ANU); L.

Kiss, University of Sydney; and A. Derekas, University of New South Wales, report that a spectrum (range 580–675 nm) of SN 2005av (cf. IAUC 8499), taken with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring Observatory on Mar. 27.76 UT, is similar to that of the type-IIn supernova 2001ir (cf. IAUC 7784).

2005 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green